### v1

@@ IV-G3 Result Composition and Presentation
The top-$k$ articles ranked by cosine similarity are retrieved along with their metadata (headline, category, and for headline matches, the full news article). The final recommendations are presented to the user as a ranked list of news articles with relevant excerpts.
@@ V Results and Discussion
This section presents a detailed analysis of the experimental results obtained from our Urdu news recommendation system. The analysis culminates in demonstrating the necessity of our proposed dual-embedding approach for handling the semantic disparity between short and long queries.
@@ V-A1 Performance on Long Queries
The superiority of mean pooling for long queries can be attributed to its ability to aggregate semantic information from all tokens in the document. Long queries, which often contain multiple contextual clues and descriptive phrases, benefit from matching against a similarly comprehensive representation of article content. This distributed semantic matching proved more effective than the focused representations from CLS pooling or the potentially noisy salient features from max pooling. Figure 19 shows the visual analysis of different pooling methods for long queries. Interestingly, in later evaluations with an expanded test set of 100 queries and retrieval depth of $k=15$, mean pooling on content embeddings for long queries achieved precision scores exceeding 90%, confirming its robustness for detailed semantic matching.
Figure 19: Visual Analysis of precision@10 for long queries using different pooling methods on content embeddings
@@ V-A2 Performance on Short Queries
Table IV presents the Precision@10 results for short queries shown in Figure 17 across the three pooling methods. CLS pooling demonstrated superior performance with an average precision of 60%, significantly outperforming mean pooling (53.07%) and max pooling (28.46%).
TABLE IV: Precision@10 for Short Queries using Different Pooling Methods on Content Embeddings
@@
The performance variance highlights a critical insight: while CLS pooling achieved the highest overall precision for short queries, even its best performance (60%) was unsatisfactory for practical recommendation scenarios. Figure 19 shows the visual analysis of different pooling methods for short queries. This indicates a fundamental limitation of using content-based embeddings for short queries. The semantic representation of a comprehensive 500-word article is too coarse and diffuse to effectively match the focused intent encapsulated in a brief query.
Figure 20: Visual Analysis of precision@10 for short queries using different pooling methods on content embeddings
The persistent under-performance on short queries necessitated the innovative solution discussed in the following Section V-B, where we introduce headline embeddings specifically designed to bridge this semantic gap.
@@ V-B Bridging the Semantic Gap: Headline Embeddings for Short Queries
The maximum precision of 60% for short queries, even with the best-performing pooling method on content embeddings, was deemed inadequate for a production-level recommendation system. This performance gap stems from a semantic granularity mismatch: short queries express focused user intent, while content embeddings represent the broad, multifaceted themes of full-length articles. This observation motivated the development of our dual-embedding approach, which seeks to align query semantics with appropriate document representations.
@@ V-B1 Hypothesis and Implementation
We hypothesized that short user queries are semantically more similar to article headlines than to full article content. Headlines, by design, are concise, information-dense summaries that capture the core topic of an article, making them better semantic matches for similarly concise queries. To test this hypothesis, we generated a new set of embeddings from the headline column of our dataset using the same pre-trained model (urduhack/roberta-urdu-small [51].). For each pooling method (mean, max, and CLS), we created separate vector databases where each headline embedding was stored alongside its corresponding full article content and category as metadata. This design ensured that when a headline matched a short query semantically, the system could retrieve and recommend the complete article to the user, not just headline.
An important distinction in processing was that headlines, being inherently short, required no chunking and were embedded in a single forward pass of the transformer model, unlike lengthy article content which required the chunking strategy described earlier.
@@ V-B2 Performance Evaluation on Short Queries
TABLE V: Performance of Headline Embeddings with Different Pooling Methods on Short Queries (Precision@10)
@@
Avg. TP 5.92 4.69 8.46
Avg. Prec. 0.5846 0.4846 0.8769
58.46% 48.46% 87.69%
Figure 21 shows the visual analysis of different pooling methods for short queries when embeddings of headline were generated.
Figure 21: Visual Analysis of precision@10 for short queries using different pooling methods on headline embeddings
@@ V-B3 Case Analysis: Precision Breakthroughs
The headline embedding approach, particularly with CLS pooling, enabled breakthroughs on queries where other methods failed catastrophically. The single-word query in Figure 22 represents the most challenging case for semantic matching.
Figure 22: Urdu short query 8
Content embeddings struggled with this query, achieving only 0.4 precision with CLS pooling. In contrast, CLS-pooled headline embeddings achieved 0.9 precision, retrieving highly relevant technology-focused headlines such as shown in Figure 23, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 23: Urdu short query 8 recommendations
Similarly query 4 as in Figure 24 exemplifies how headline embeddings better capture domain-specific semantics.
Figure 24: Urdu short query 4
With content embeddings, precision was only 0.5, but headline embeddings with CLS pooling achieved 0.8 precision, correctly retrieving election-related headlines as shown in Figure 25, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 25: Urdu short query 4 recommendations
Query 5 as shown in Figure 26 demonstrates the improved specificity of headline embeddings.
Figure 26: Urdu short query 5
Previous methods retrieved general entertainment news from Bollywood and Hollywood due to the broad semantic representation in content embeddings. However, CLS-pooled headline embeddings correctly captured the nationality modifier ”Pakistani”, retrieving exclusively Pakistan-centric entertainment headlines achieving 0.9 precision as illustrated in Figure 27, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 27: Urdu short query 5 recommendations
@@
TABLE VI: Comparative Precision@10 for Short Queries: Content vs. Headline Embeddings
@@
Figure 28 shows the line chart comparison of all pooling methods using both content and headline embeddings.
@@ V-B4 Implications for Pooling Strategy Selection
This optimized configuration leverages the strengths of each approach: mean pooling’s comprehensive aggregation for matching long, detailed queries against rich article content, yielding optimal results (74% precision) and CLS pooling’s focused semantic capture for matching short, intent-driven queries against concise headlines. The 87.69% precision achieved for short queries represents not just a quantitative improvement, but a qualitative transformation in the system’s ability to understand and respond to focused user intent.
@@ V-C Dimensionality Reduction for Efficient Retrieval
While our dual-embedding approach established an optimal framework for semantic matching, operational efficiency remained a critical concern. With our dataset scaling to 112,000 records, performing nearest-neighbor searches in the original 768-dimensional space incurred significant computational and temporal costs. To enhance the scalability and real-time responsiveness of our recommendation system without substantially compromising retrieval quality, we investigated dimensionality reduction (DR) techniques. The objective was to find lower-dimensional representations that preserved the essential semantic relationships within both our content and headline embedding spaces.
@@ V-C1 Dimensionality reduction for content embeddings
This metric provides a normalized measure of set similarity that accounts for both true positives and the union of retrieved items, offering a robust assessment of how well the reduced space preserves the retrieval characteristics of the original embedding space.
@@
Table VII presents the comparative performance of the three DR techniques on content embeddings. PCA demonstrated decisive superiority, achieving a mean overlap score of 74.92% and a mean Jaccard Index of 0.572.
@@
Mean 74.92% 32.46% 42.31% 37.46 16.23 21.15 0.572 0.187 0.273
Figures 29 and 30 present the evaluation results of dimensionality reduction techniques applied to content embeddings evaluated on long test queries. Figure 29 shows a horizontal bar graph of the top-50 recommendation overlap between 64-dimensional reduced embeddings and the 768-dimensional ground truth, while Figure 30 displays the corresponding average Jaccard index distribution across techniques.
Figure 29: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth: overlap in top-50 recommendation across DR techniques for content vector DB
Figure 30: Comparative analysis of Jaccard Index distributions for evaluating DR technique fidelity to original semantic space
The results reveal a clear hierarchy: PCA significantly outperformed both UMAP and the Autoencoder. The linear projection of PCA proved exceptionally effective at preserving the semantic neighborhood structure crucial for ranking a large set of similar articles. While the Autoencoder showed reasonable performance on several queries, its inconsistency was notable (e.g., Q5 with 0% overlap), suggesting potential instability or overfitting during training. UMAP, despite its strength in visualizing complex manifolds, performed the poorest in this retrieval-focused task, indicating that its non-linear transformations may distort the global similarity relationships needed for effective nearest-neighbor search.
@@ V-C2 Dimensionality reduction for headline embeddings
Following the conclusive results for content embeddings, we systematically applied the same evaluation framework to the headline embedding space, the critical component of our short query pathway. The identical three techniques (PCA, UMAP, and Autoencoder) were employed to reduce the 768-dimensional CLS-pooled headline embeddings to 64 dimensions, as CLS pooling was established in Section  V-B4 as the optimal strategy for generating semantic representations of headlines. The remaining configurations are same as were adopted in case of DR analysis for content embeddings, dicussed in Section  V-C1. The evaluation utilized the same 13 short queries as shown in Figure 17 that defined the short query category, with the original 768D headline embeddings serving as the ground truth. Retrieval depth was set to $K=50$ to provide a rigorous test of each technique’s ability to preserve ranking fidelity across a broad set of similar items.
@@ V-C3 Results and Comparative Analysis
The performance metrics for headline embeddings are presented in Table VIII. PCA once again demonstrated clear superiority, achieving a mean overlap score of 65.10% and a mean Jaccard Index of 0.471.
@@
Mean 65.10% 17.50% 50.46% 32.55 8.75 25.23 0.471 0.090 0.324
Figures 31 and 32 present the evaluation results of dimensionality reduction techniques applied to headline embeddings evaluated on short test queries defined in Figure 17 . Figure 31 shows a horizontal bar graph of the top-50 recommendation overlap between 64-dimensional ‘reduced embeddings and the 768-dimensional ground truth, while Figure 32 displays the corresponding average Jaccard index distribution across techniques.
Figure 31: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth for headline vector DB: overlap in top-50 across DR techniques
Figure 32: Comparative analysis of Jaccard Index distributions for evaluating DR technique fidelity to original semantic space for headline vector DB
@@ V-D Determining Optimal Embedding Dimension
Following the selection of Principal Component Analysis (PCA) as the superior dimensionality reduction technique, we sought to identify the optimal reduced dimension size for each pathway of our dual-embedding architecture. The primary objective was to maximize the preservation of semantic fidelity relative to the original 768-dimensional space while achieving a practical dimensionality for efficient storage and retrieval. This trade-off is inherent: lower dimensions may lead to greater information loss, while higher dimensions offer diminishing returns in semantic preservation at the cost of increased computational and storage overhead.
To systematically explore this trade-off, we evaluated three candidate dimensions, 64, 128, and 256 selected as powers of two for alignment with computational best practices in vector processing. The evaluation was conducted separately for content embeddings (serving long queries) and headline embeddings (serving short queries) to identify dimension-specific optima tailored to each semantic space.
@@ V-D1 Optimal Dimension for Long Queries: Content Embeddings
In the preceding dimensionality reduction analysis, a fixed target dimension of 64D was used for all candidate techniques (PCA, UMAP, Autoencoder) to ensure a fair and consistent comparison of their ability to preserve semantic information. However, to optimize the system for practical deployment, it is essential to determine the specific reduced dimension that best balances retrieval speed with semantic fidelity for each embedding pathway. The trade-off is inherent: lower dimensions yield faster computations but risk greater information loss, while higher dimensions preserve more semantics at increased computational cost.
[…]
It is important to clarify the operational definition of a ”long query” in our system. Beyond traditional short search- bar queries, we consider the full text of an article a user is actively reading as a long query. This scenario is common in content recommendation systems, where the article being viewed serves as the contextual seed for generating related recommendations. In this case, the article’s entire content acts as the query, triggering the long-query pipeline. Empirical observation indicates that recommendation precision generally improves with query length, as richer context allows for more accurate semantic matching.
@@
The performance of PCA-reduced content embeddings across dimensions and query lengths is detailed in Table IX. The values represent averages computed across the 100 test queries for each length interval. Figures  33 and 34 shows grouped bars and heatmap analysis of statistics in table Table IX.
TABLE IX: Performance of PCA-reduced content embeddings across dimensions and query lengths
Query Length Average Overlap with 768D for 100 Test Queries(%)
(Characters) 64D 128D 256D
150 72.8 76.4 76.7
200 74.6 76.5 76.3
250 73.9 76.2 75.9
300 73.1 75.9 75.5
350 72.7 74.4 74.3
Mean 73.4 75.9 75.8
@@
Figure 34: Heatmap analysis of overlap performance comparison across PCA dimensions and query lengths for content embeddings

### v2

@@ IV-G3 Result Composition and Presentation
The top-$k$ highest ranking articles by cosine similarity are returned together with their metadata (headline, category, and in the case of headline matches, the entire news article). The user is given the final recommendations in a list of relevant news articles with excerpts ranked.
@@ V Results and Discussion
This section presents a systematic experimental evaluation of our proposed Urdu news recommendation system. We begin in Section V-A by comparing mean, max, and CLS pooling strategies to identify the optimal method for generating article embeddings, revealing distinct performance patterns for short queries (Section V-A2) versus long queries (Section V-A1). Based on the observed under-performance on short queries, we introduce and evaluate headline embeddings as a specialized solution in Section V-B, demonstrating their superiority (Section V-B2) and analyzing breakthrough cases (Section V-B3) that validate our core hypothesis. To address scalability with 112k articles, Section V-C evaluates PCA, UMAP, and Autoencoder for dimensionality reduction, separately analyzing performance on content embeddings (Section V-C1) and headline embeddings (Section V-C2). Section V-D determines the optimal reduced dimensions for each pathway i.e., for content embeddings (Section V-D1) and for headline embeddings (Section V-D2). Finally, Section V-E presents the complete dual-pathway configuration and evaluates its performance on 100 test queries per pathway (Section V-E1), followed by comparative analysis with a baseline system (Section V-F).
@@ V-A1 Performance on Long Queries
The superiority of mean pooling for long queries can be attributed to its ability to aggregate semantic information from all tokens in the document. Long queries, which often contain multiple contextual clues and descriptive phrases, benefit from matching against a similarly comprehensive representation of article content. This distributed semantic matching proved more effective than the focused representations from CLS pooling or the potentially noisy salient features from max pooling. Interestingly, in later evaluations with an expanded test set of 100 queries and retrieval depth of $k=15$, mean pooling on content embeddings for long queries achieved precision scores exceeding 90%, confirming its robustness for detailed semantic matching.
@@ V-A2 Performance on Short Queries
The Precision@10 results for short queries displayed in Figure 13 is provided in Table IV across the three pooling methods. CLS pooling exhibited better performance with an average precision of 60%, outperforming mean pooling (53.07%) and max pooling (28.46%).
TABLE IV: Performance of content embeddings using different pooling methods on short queries (Precision@10)
@@
The performance variance illustrates an important point: although CLS pooling demonstrated the best overall precision with short queries, its best performance (60%) was not satisfactory in practice, when it comes to recommendations. This means that there is an inherent weakness in applying content-based embeddings to short queries. The semantic expression of a detailed article of 500 words is too broad to effectively correspond to the narrow intent that is captured in a short query.
This continued under-performance on short queries led to the innovative solution presented in the next Section V-B, where we present headline embeddings that are specifically created to address this semantic gap.
@@ V-B Bridging the Semantic Gap: Headline Embeddings for Short Queries
The highest precision of 60% on short queries, even using the most effective pooling method on content embeddings, was insufficient to a production-scale recommendation system. This gap in performance is caused by a mismatch in semantic granularity: short queries reflect narrow user intent, and content embeddings reflect the multidimensional themes of full-length articles. This observation inspired the design of our dual-embedding strategy, which aims to match query semantics with relevant document representations.
@@ V-B1 Hypothesis and Implementation
Our hypothesis was that short user queries were semantically closer to article headlines than to the content of full articles. Headlines are, by definition, short, information-dense summaries of the essence of an article, and are therefore more semantically aligned to short queries. To test this hypothesis, we produced a new set of embeddings of the headline column of our dataset using the same pre-trained model (urduhack/roberta-urdu-small [51]). For all pooling techniques (mean, max, and CLS), we prepared individual pooling vector databases in which every headline embedding was stored with its respective full article text and category as metadata. This architecture guaranteed that in case a headline corresponded with a brief query in semantic terms, the system would be able to fetch and suggest the entire article to the user, not just headline.
A key difference in processing was that headlines were short enough to not need chunking and could be processed in single forward pass of the transformer model whereas long article text needed chunking outlined earlier.
@@ V-B2 Performance Evaluation on Short Queries
TABLE V: Performance of headline embeddings using different pooling methods on short queries (Precision@10)
@@
Average 0.5846 0.4846 0.8769
Mean Precision 58.46% 48.46% 87.69%
@@ V-B3 Case Analysis: Precision Breakthroughs
The headline embedding approach, particularly with CLS pooling, enabled breakthroughs on queries where other methods failed catastrophically. A few such examples of precision breakthroughs are discussed below. The single-word query in Figure 15 represents the most challenging case for semantic matching. Content embeddings struggled with this query, achieving only 40% precision with CLS pooling. In contrast, CLS-pooled headline embeddings achieved 90% precision corresponding to 9 TPs out of 10, retrieving highly relevant technology-focused headlines such as shown in Figure 15, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 15: Urdu short query 8 recommendations example 1
Similarly another short query as in Figure 16 exemplifies how headline embeddings better capture domain-specific semantics. With content embeddings, precision was only 50%, but headline embeddings with CLS pooling achieved 80% precision meaning there were 8 TPs, correctly retrieving election-related headlines as shown in Figure 16, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 16: Urdu short query recommendations example 2
Another query as shown in Figure 17 demonstrates the improved specificity of headline embeddings. Previous methods retrieved general entertainment news from Bollywood and Hollywood due to the broad semantic representation in content embeddings. However, CLS-pooled headline embeddings correctly captured the nationality modifier ”Pakistani”, retrieving exclusively Pakistan-centric entertainment headlines achieving 90% precision as illustrated in Figure 17, where 1 indicates relevant query, and 0 indicates irrelevant.
Figure 17: Urdu short query recommendations example 3
@@
TABLE VI: Comparative Precision@10 for short queries: Content vs. headline embeddings using CLS pooling method
@@
Figure 18 shows the line chart comparison of all pooling methods using both content column and headline column embeddings.
@@ V-B4 Implications for Pooling Strategy Selection
This is an optimized design that exploits the advantages of both designs: comprehensive aggregation of mean pooling to match long and detailed queries to rich article content achieving the best results (74% precision) and focused semantic capture of CLS pooling to match short and purposeful queries to brief headlines. The 87.69% precision on short queries is not only a quantitative change, but a qualitative shift in how the system comprehends and acts on focused user intent.
@@ V-C Dimensionality Reduction for Efficient Retrieval
Although our dual-embedding methodology created optimal structure for semantic matching, there was still a concern on operational efficiency. With extremely large corpus size, nearest-neighbor searches in the original 768-dimensional space were very expensive in terms of computational and time costs. To improve the scalability of our recommendation system and its real-time responsiveness without significantly reducing the quality of retrieval, we explored the methods of dimensionality reduction (DR). It attempted to identify lower-dimensional representations that could maintain the key semantic associations within both content and headline embedding spaces.
@@ V-C1 Dimensionality reduction for content embeddings
This measure gives a normalized quantification of set similarity taking into consideration true positives and the union of retrieved items, which is a strong measure of the extent to which the reduced space maintains the retrieval properties of the original embedding space.
@@
Table VII presents the comparative performance of the three DR techniques on content embeddings. PCA demonstrated decisive superiority, achieving a mean overlap score of 74.92% and a mean Jaccard Index of 0.572 as shown in Figures 19 and 20.
@@
Figures 19 and 20 present the evaluation results of dimensionality reduction techniques applied to content embeddings evaluated on long test queries. Figure 19 shows a horizontal bar graph of the average of top-50 recommendation overlap between 64-dimensional reduced embeddings and the 768-dimensional ground truth, while Figure 20 displays the corresponding average Jaccard index distribution across techniques.
Figure 19: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth: average overlap in top-50 recommendation across DR techniques for content vector DB
Figure 20: Comparison of average Jaccard Index distributions to analyze fidelity of DR technique to original semantic space for content vector DB
The results reveal a clear hierarchy: PCA significantly outperformed both UMAP and the Autoencoder. The PCA linear projection was extremely successful in maintaining the semantic neighborhood structure that is important in ranking a large number of similar articles. Although the Autoencoder was performing reasonably on multiple queries, it was quite inconsistent (e.g. Q5 with 0% overlap). UMAP, although it is strong at visualizing complex manifolds, had the worst performance in this retrieval-based task, suggesting that its non-linear transformations can lead to distortion of the global similarity relationships critical to a successful nearest-neighbor search.
@@ V-C2 Dimensionality reduction for headline embeddings
After the final findings on the content embeddings, we applied the same evaluation systematically to headline embeddings space which is the crucial component of our short query pathway. The same three methods (PCA, UMAP, and Autoencoder) were used to further reduce the 768-dimensional CLS-pooled headline embeddings to 64 dimensions since CLS pooling was determined in Section  V-B4 as the best approach to provide semantic representations of headlines. The remaining configurations are same as were adopted in case of DR analysis for content embeddings, discussed in Section  V-C1. The evaluation utilized the same 13 short queries as shown in Figure 13 that defined the short query category, with the original 768D headline embeddings serving as the ground truth. Retrieval depth was set to $K=50$ to provide a rigorous test of each technique’s ability to preserve ranking fidelity across a broad set of similar items.
@@ V-C3 Results and Comparative Analysis
The performance metrics for headline embeddings are presented in Table VIII. PCA once again demonstrated clear superiority, achieving a mean overlap score of 65.10% and a mean Jaccard Index of 0.471 illustrated in Figures 21 and 22.
@@
Figures 21 and 22 present the evaluation results of dimensionality reduction techniques applied to headline embeddings evaluated on short test queries defined in Figure 13 . Figure 21 shows a horizontal bar graph of the average top-50 recommendation overlap between 64-dimensional ‘reduced embeddings and the 768-dimensional ground truth, while Figure 22 displays the corresponding average Jaccard index distribution across techniques.
Figure 21: Fidelity of 64-dimensional embeddings to the 768-dimensional ground truth for headline vector DB: average overlap in top-50 across DR techniques
Figure 22: Comparison of average Jaccard Index distributions to analyze fidelity of DR technique to original semantic space for headline vector DB
@@ V-D Determining Optimal Embedding Dimension
Following the selection of Principal Component Analysis (PCA) as the superior dimensionality reduction technique, we sought to identify the optimal reduced dimension size for each pathway of our dual-embedding architecture. The primary objective was to maximize the preservation of semantic fidelity relative to the original 768-dimensional space while achieving a practical dimensionality for efficient storage and retrieval. This is an inevitable trade-off: reducing dimensions can result in an increased loss of information, whereas increasing dimensions is a diminishing game in terms of semantic preservation, and raising the cost of computation and storage.
To comprehensively investigate this trade-off, we considered three candidate dimensions, 64, 128, and 256 chosen as powers of 2 to match to computational best practices in the area of vector processing. The analysis was done on both content embeddings (serving long queries) and headline embeddings (serving short queries) to determine dimension-specialized optima in each semantic space.
@@ V-D1 Optimal Dimension for Long Queries: Content Embeddings
In the above dimensionality reduction analysis, all candidate techniques (PCA, UMAP, Autoencoder) used a target dimension of 64D to guarantee a fair and balanced comparison for regular assessment of their capacity to maintain semantic information. But to make system practical, it is necessary to identify the particular reduced dimension that optimizes speed of retrieval and semantic fidelity to every embedding pathway. This trade-off is implicit: lower dimensions imply faster computations but often lead to higher loss of information, while higher dimensions retain semantics at higher computation costs.
In our dual-embedding architecture query routing is controlled by a character-level cutoff, denoted by $\theta$, as established in Section IV-A2. The threshold $\theta$ = 150 characters was determined empirically based on the maximum headline length in our corpus, ensuring that queries semantically comparable to headlines are routed to the headline embedding pathway. This threshold selection principle is generalizable: $\theta$ should be informed by the characteristic length of titles, headings, or product titles in a given recommender system’s content corpus.
It is important to clarify the operational definition of a ”long query” in our system. In addition to conventional short search- bar queries, we regard the entire text of an article a user is reading actively as a long query. Such a situation is typical in content recommendation systems, in which the article currently viewed acts as the contextual seed to produce related recommendations. In this instance, the whole content of the article acts as query and activates the long-query pipeline. Empirical evidence shows that the relevance of recommendations tends to rise with the length of query, with a more informative context permitting it to better match semantically.
@@
The performance of PCA-reduced content embeddings across dimensions and query lengths is detailed in Figures  23 through grouped bars. The values represent averages computed across the 100 test queries for each length interval.